A rectangular glass block (n=1.52) is placed inside a clean water contained in a basin. A layer of oil (n=1.44) floats on the water surface. A ray of light from air strikes the oil surface at an angle of incidence of 28° . Calculate the angle of refraction at the water-glass interface. (refractive index of air =1.00)
1
Expert's answer
2019-12-16T11:00:39-0500

air-oil

"\frac{sin(\alpha_1)}{sin(\alpha_2)}=\frac{n_o}{n_a}"

oil-water

"\frac{sin(\alpha_2)}{sin(\alpha_3)}=\frac{n_w}{n_o}"

water-glass

"\frac{sin(\alpha_3)}{sin(\alpha_4)}=\frac{n_g}{n_w}"

combine this equation


"sin(\alpha_4)=\frac{n_w}{n_g}\cdot \frac{n_o}{n_w}\cdot \frac{n_a}{n_o}\cdot sin(\alpha_1)""\alpha_4=arcsin(\frac{n_a}{n_g}\cdot sin(\alpha_1))"

calculating

"\alpha_4=arcsin(\frac{1}{1.52}\cdot sin(28))=18^o"